OpenStudy (anonymous):

express the following in simplest radical form. square root of 50

OpenStudy (anonymous):

\(50=25\times 2\) and so \[\sqrt{50}=\sqrt{25\times 2}=\sqrt{25}\times \sqrt{2}=5\sqrt2\]